### v1

@@ 4.1 A Constant Fraction of Points Inside and Outside Geodesic Disks
Now consider some fixed bisector $b(p,q)$. As in the proof of Theorem 3, the edges of the order-$1$ through order-$k$ diagrams on $b(p,q)$ form a number of connected components (which is strictly less than the total number of edges of all the diagrams). Let $\lambda_{k}$ be the number of connected components formed by all the bisectors of the order-$1$ through order-$k$ diagrams. Analogously, the edges of the order-$(n-k)$ through order-$(n-1)$ diagrams on $b(p,q)$ form another number of connected components. Let $\overline{\lambda}_{n-k}$ be the number of these connected components formed by all the bisectors of the order-$(n-k)$ through order-$(n-1)$ diagrams. The goal is to find the largest value of $k$ such that $\lambda_{k}+\overline{\lambda}_{n-k}<\binom{n}{2}$. When $\lambda_{k}+\overline{\lambda}_{n-k}<\binom{n}{2}$, we note that there exists a bisector $b(p,q)$ that does not have any edges of the order-$1$ through order-$k$ diagrams, and of the order-$(n-k)$ through order-$(n-1)$ diagrams. Therefore, every geodesic disk with $p$ and $q$ on its boundary has at least $k$ points, both inside and outside the disks.
Since $\overline{\lambda}_{n-k}=\lambda_{n-1}-\lambda_{n-k-1}$, one can follow the reasoning from Theorem 3 to show that $\lambda_{j}=\sum_{i=1}^{j}(-1)^{j-i}e_{i}+\left(S_{j-1}+S_{j-3}+\ldots+S_{2\mbox{-}or\mbox{-}1}\right)$, for $j\in\{1,\ldots,n-1\}$, where $e_{j}$ is the number of edges in an order-$j$ diagram and $S_{j}$ is the number of edges of the order-$j$ diagram that intersect the boundary. Since an order-$j$ geodesic Voronoi diagram is an order-$j$ abstract Voronoi diagram, we have that $e_{j}=6jn-3j^{2}-3n-3\sum_{i=1}^{j-1}S_{i}-S_{j}$, see [8, Lemmas $15$ and $16$]. Then we get Equation (2), for $j\in\{2,\ldots,n-1\}$. In the final calculation we distinguish the case of $k$ being even from that of $k$ being odd. However, in both cases we have that
$\lambda_{k}+\lambda_{n-1}-\lambda_{n-k-1}=3kn-\sum_{i=1}^{k}S_{i}-\sum_{i=n-k}^{n-1}S_{i}=3kn-\sum_{i=1}^{k}S_{i}-\sum_{i=1}^{n-1}S_{i}+\sum_{i=1}^{n-k-1}S_{i}.$
If we now use the fact that $j(j+1)\leq\sum_{i=1}^{j}S_{i}\leq j(2n-j-1)$, see [8, Lemma $11$], we get that
$\lambda_{k}+\lambda_{n-1}-\lambda_{n-k-1}\leq 3kn-2k(k+1)=3kn-2k^{2}-2k,$
which is a quadratic polynomial in $k$. We note that $\lambda_{k}+\overline{\lambda}_{n-k}<\binom{n}{2}$ whenever $k<\frac{n}{3+\sqrt{5}}$. Therefore, the same arguments at the end of the proof of Theorem 3 lead to $\Pi^{in-out}(n)\geq\left\lceil\frac{n}{3+\sqrt{5}}\right\rceil+1$. ∎
@@ 4.2 Sets in geodesically convex position
In the Euclidean setting, Hayward et al. [14] proved that when points are in convex position, for every disk containing two points on the boundary, there always exists a pair of points such that any disk containing that pair contains at least $\left\lceil\frac{n}{3}\right\rceil+1$. We generalize that result to the geodesic setting by showing that $\overline{\Pi}(n)\geq\left\lceil\frac{n}{3}\right\rceil+1$. The geodesic case poses its unique challenges, for example, it is not always possible to find an enclosing geodesic disk with three points on the boundary. However, when no such enclosing geodesic disk exists, we show that there exists a pair of points such that every geodesic disk through them is an enclosing geodesic disk.
@@
Define the notation $D(c)$ to be the disk $D(c,\max_{p\in S}|g(c,p)|)$ which is the geodesic disk centered at an arbitrary point $c\in P$, and having as radius the maximum geodesic distance from $c$ to any other point $p\in S$. By construction, this is an enclosing geodesic disk. Consider an arbitrary disk $D(x)$ centered on a point $x\in S$ with a point $y\in S$ on the boundary of $D(x)$. Either $D(x)$ has a second point $z$ on its boundary or only has $y$ on its boundary. If the latter is true, then we show how to obtain an enclosing disk with a second point on its boundary. Consider the disk $D(x)$ as the center $x$ moves along $g(x,y)$ towards $y$. Since the radius is shrinking, eventually, a second point $z\in S$ will reach the boundary of $D(x)$. Note that $D(x)$ always remains enclosing.
Thus, $D(x)$ is an enclosing disk with $y$ and $z$ on its boundary. We now show that every geodesic disk through $y$ and $z$ is enclosing. Let $w$ be any point on $b(y,z)$. $D(w)$ has $y$ and $z$ on its boundary. Moreover, $D(w)$ must be enclosing. If $D(w)$ is not enclosing, then we contradict the fact that there is no enclosing geodesic disk through three points. The lemma follows. ∎
We now generalize the result from Hayward et al. [14] for point sets in convex position to the geodesic setting. In particular, when $S$ is a set of points in $P$ that is in geodesically convex position, then there is a pair of points such that any geodesic disk through them contains at least $\left\lceil\frac{n}{3}\right\rceil+1$ points of $S$.
@@
For any set $S$ of $n$ points in geodesically convex position in a polygon, $\overline{\Pi}(n)\geq\left\lceil\frac{n}{3}\right\rceil+1$.
@@
Otherwise, there is an enclosing disk $D$ that has three points on its boundary, namely $x,y,z$ in clockwise order along the boundary of $D$. Let $f(\triangle(x,y,z))$ be the maximum among the number of points from $S$ contained in each of the three regions $D\cap P^{+}(x,y),D\cap P^{+}(y,z)$ and $D\cap P^{+}(z,x)$. Among all enclosing disks with 3 points on the boundary, consider disk $D_{1}$ with points $u,v,w\in S$ on its boundary which minimizes $f(\triangle(u,v,w))$. There must be at least one region among $D_{1}\cap P^{+}(u,v),D_{1}\cap P^{+}(v,w)$ and $D_{1}\cap P^{+}(w,u)$ whose interior has at least $\left\lceil\frac{n-3}{3}\right\rceil$ points. Assume, without loss of generality, that $D_{1}\cap P^{+}(u,v)$ does.
@@
If the interior of $D_{1}\cap P^{-}(u,v)$ contains fewer than $\left\lceil\frac{n-3}{3}\right\rceil$ points, then we continuously change $D_{1}$ by the disks through $u,v$ obtained when moving the center $c$ of $D_{1}$ along the part of the bisector $b(u,v)$ to the right of $\ell(u,c)$, as in Figure 5. While doing this, all those disks will keep being enclosing disks, by Lemma 1. As we move $c$, we may arrive at one of two situations: Either the center $c$ hits the boundary of the polygon, or the boundary of the current disk, $D_{2}$, hits a point $w_{1}\in D_{1}\cap P^{+}(u,v)$. We show, by contradiction, that the latter case cannot occur. If it occurred, the geodesic triangle $\triangle(u,w_{1},v)$ divides $D_{2}$ into three regions $D_{2}\cap P^{+}(u,w_{1})$, $D_{2}\cap P^{+}(w_{1},v)$ and $D_{2}\cap P^{+}(v,u)$. Note that $D_{2}\cap P^{+}(v,u)=D_{2}\cap P^{-}(u,v)$ contains fewer than $\left\lceil\frac{n-3}{3}\right\rceil$ points of $S$, since it is contained in $D_{1}\cap P^{-}(u,v)$. Also, $D_{2}\cap P^{+}(u,w_{1})$ and $D_{2}\cap P^{+}(w_{1},v)$ both contain fewer points than $D_{1}\cap P^{+}(u,v)$. Thus, $f(\triangle(u,w_{1},v))<f(\triangle(u,v,w))$, and contradicting the minimality of $f(\triangle(u,v,w))$.
@@ 5 Diametral geodesic disk
Let $D$ be an enclosing geodesic disk with three points on its boundary. If such a disk does not exist, by Lemma 8, there exist two points $x,y\in S$ such that every geodesic disk through them is a geodesic enclosing disk. In particular, the disk $D\left(c_{xy},\frac{|g(x,y)|}{2}\right)$ is an enclosing diametral disk with $x$ and $y$ as its diametral endpoints.
Otherwise, let $x,y,z\in S$ be the three points on the boundary of $D$, in clockwise order. If the center of the disk is inside the geodesic triangle $\triangle(x,y,z)$, we are done. Otherwise, the center is inside one of the regions $D\cap P^{+}(x,y),D\cap P^{+}(y,z)$ or $D\cap P^{+}(z,x)$. Assume, without loss of generality, it is in $D\cap P^{+}(x,y)$. Then we move the center $c$ of $D$ along $b(x,y)$ towards $c_{xy}$. If the center $c$ reaches $c_{xy}$, then $D\left(c_{xy},\frac{|g(x,y)|}{2}\right)$ is a diametral enclosing geodesic disk with $x$ and $y$ as its diametral endpoints. If the disk hits a third point $w\in S$, and the center of the disk is inside $\triangle(x,y,w)$ , we are done. Otherwise, we repeat the process with the points $x,y,w$ as the new points $x,y,z$ until the center of the disk through the three points on the boundary of the disk lies inside the geodesic triangle formed by the three points on the boundary. This process ends because the distance to the center from the closest shortest path is decreasing at every iteration. ∎
@@
For any set $S$ of $n$ points in a polygon, $\Pi^{diam}(n)=\overline{\Pi}^{diam}(n)\geq\left\lceil\frac{n}{3}\right\rceil+1$.
@@
For case (i), let $a,b$ and $c$ be the three points on the boundary of the geodesic disk. We will prove that any point $x\in S$ is contained in at least one of the geodesic disks $D\left(c_{ab},\frac{|g(a,b)|}{2}\right)$, $D\left(c_{bc},\frac{|g(b,c)|}{2}\right)$ and $D\left(c_{ca},\frac{|g(c,a)|}{2}\right)$.
First, consider the case where $x$ is outside the geodesic triangle $\triangle(a,b,c)$. Then, $x$ belongs to one of the regions $D\cap P^{+}(a,b),D\cap P^{+}(b,c)$, and $D\cap P^{+}(a,c)$. Without loss of generality, assume $x\in D\cap P^{+}(a,b)$. By Lemma 9, the center of the geodesic disk is inside $\triangle(a,b,c)$, so it does not belong to $D\cap P^{+}(a,b)$. Then, by Lemma 1, $x$ is contained in the diametral geodesic disk $D\left(c_{ab},\frac{|g(a,b)|}{2}\right)$.
Now, we prove the case where $x\in\triangle(a,b,c)$. Consider the geodesic paths $g(a,x),g(b,x)$ and $g(c,x)$. One of the angles $\angle axb,\angle bxc$ or $\angle axc$ is greater than or equal to $\frac{2\pi}{3}$, and the other two angles are smaller than or equal to $\frac{\pi}{3}$. Assume, without loss of generality, that $\angle bxc\geq\frac{2\pi}{3}$, see blue angle in Figure 6. Let $o$ be the midpoint of the shortest path $g(b,c)$. We also know that one of the angles $\angle bxo$ or $\angle oxc$ is greater than or equal to $\frac{\pi}{3}$. Assume, without loss of generality, that it is the angle $\angle oxc$. When $o$ is not visible from $x$, the angle $\angle oxc$ is equal to the angle $\angle bxc\geq\frac{2\pi}{3}>\frac{\pi}{2}$. Hence, by [20, Corollary $2$], $|g(x,o)|<|g(o,c)|=|g(o,b)|$, and $x$ is contained in the geodesic disk $D\left(c_{bc},\frac{|g(b,c)|}{2}\right)$. In the case of $o$ being visible from $x$, by Lemma 4, $|g(o,c)|\geq\min\{|g(c,x)|,|g(x,o)|\}$ and, by Lemma 5, $|g(x,o)|\leq\max\{|g(c,x)|,|g(o,c)|\}$, hence $|g(x,o)|\leq|g(o,c)|=|g(o,b)|$, and $x$ is contained in the geodesic disk $D\left(c_{bc},\frac{|g(b,c)|}{2}\right)$.
@@
Then, the union of the geodesic disks $D\left(c_{ab},\frac{|g(a,b)|}{2}\right),D\left(c_{bc},\frac{|g(b,c)|}{2}\right)$ and $D\left(c_{ca},\frac{|g(c,a)|}{2}\right)$ covers all the points in $P$. Since the points $a,b$ and $c$ are counted twice,
$\Bigg{\lvert}D\left(c_{ab},\frac{|g(a,b)|}{2}\right)\cap P\Bigg{\rvert}+\Bigg{\lvert}D\left(c_{bc},\frac{|g(b,c)|}{2}\right)\cap P\Bigg{\rvert}+\Bigg{\lvert}D\left(c_{ca},\frac{|g(c,a)|}{2}\right)\cap P\Bigg{\rvert}>n+3.$

### v2

@@ 4.2 A Constant Fraction of Points Inside and Outside Geodesic Disks
Now consider some fixed bisector $b(p,q)$, and let $b_{k}(p,q)$ be the union of the edges of the order-$1$ through order-$k$ diagrams on $b(p,q)$. As in the proof of Theorem 3, the edges of $b_{k}(p,q)$ form a number of connected components (which is strictly less than the total number of edges of all the diagrams). Let $\lambda_{k}$ be the sum of the number of connected components formed by the edges of $b_{k}(p,q)$ over all pairs $p,q\in S$. Analogously, let $\overline{b}_{k}(p,q)$ be the union of the edges of the order-(nk)(n-k) through order-(n1)(n-1) diagrams on $b(p,q)$. The edges of $\overline{b}_{k}(p,q)$ form a number of connected components. Let $\overline{\lambda}_{n-k}$ be the sum of the number of these connected components formed by the edges of $\overline{b}_{k}(p,q)$ over all pairs $p,q\in S$. The goal is to find the largest value of $k$ such that $\lambda_{k}+\overline{\lambda}_{n-k}<\binom{n}{2}$. When $\lambda_{k}+\overline{\lambda}_{n-k}<\binom{n}{2}$, we note that there exists a bisector $b(p,q)$ that does not have any edges of the order-$1$ through order-$k$ diagrams, and of the order-(nk)(n-k) through order-(n1)(n-1) diagrams. Therefore, every geodesic disk with $p$ and $q$ on its boundary has at least $k$ points, both inside and outside the disks.
We already have an upper bound on $\lambda_{k}$ from the proof of Theorem 3, so we will now provide a similar upper bound on $\overline{\lambda}_{n-k}$. As in the proof of Theorem 3, we use the term endpoint to denote Voronoi vertices of the order-$k$ diagrams, not on the polygon boundary. We define $\overline{\phi}_{n-k}$ to be the number of unpaired endpoints after putting down all the edges of the order-(n1)(n-1) through the order-(nk)(n-k) geodesic Voronoi diagrams. One can follow similar reasoning to the proof of Theorem 3 to arrive at the formula
$\overline{\lambda}_{n-k}=\sum_{j=1}^{k}e_{n-j}-\sum_{j=1}^{k-1}\overline{\phi}_{n-j}.$
We also have $\overline{\phi}_{n-k}+\overline{\phi}_{n-k+1}=2e_{n-k}-S_{n-k}$. Note that $\overline{\phi}_{n-1}=2e_{n-1}-S_{n-1}$ since if we only put down edges of the order-(n1)(n-1) diagram, every endpoint will be unpaired, and $2e_{n-1}-S_{n-1}$ is the number of endpoints in the order-(n1)(n-1) diagram. Finally, recall that Bohler et al. [8, Lemma 11] proved $j(j+1)\leq\sum_{i=1}^{j}S_{i}\leq j(2n-j-1)$.
As in the proof of Theorem 3, we consider the cases where $k$ is even or odd separately. When $k$ is even, we have
$\overline{\lambda}_{n-k}$ $=\sum_{j=1}^{k}e_{n-j}-\sum_{j=1}^{k-1}\overline{\phi}_{n-j}$
$=\sum_{j=1}^{k}e_{n-j}-[(\overline{\phi}_{n-k+1}+\overline{\phi}_{n-k+2})+(\overline{\phi}_{n-k+3}+\overline{\phi}_{n-k+4})+\ldots+(\overline{\phi}_{n-3}+\overline{\phi}_{n-2})+\overline{\phi}_{n-1}]$
$=\sum_{j=1}^{k}e_{n-j}-[(2e_{n-k+1}-S_{n-k+1})+(2e_{n-k+3}-S_{n-k+3})+\ldots+(2e_{n-1}-S_{n-1})]$
$=\sum_{j=1}^{k}(-1)^{k-j}e_{n-j}+(S_{n-k+1}+S_{n-k+3}+\ldots+S_{n-1})$
$=-\sum_{j=1}^{k/2}(e_{n-2j+1}-e_{n-2j})+\sum_{j=1}^{k/2}S_{n-2j+1}$
$=-\sum_{j=1}^{k/2}(6n-6(n-2j+1)+3-S_{n-2j+1}-2S_{n-2j})+\sum_{j=1}^{k/2}S_{n-2j+1}$
$=-\sum_{j=1}^{k/2}(6n-6(n-2j+1)+3)+2\sum_{j=1}^{k/2}(S_{n-2j+1}+S_{n-2j})$
$=-\sum_{j=1}^{k/2}(12j-3)+2\sum_{j=1}^{k}S_{n-j}$
$=-\frac{3}{2}k(k+1)+2\bigg(\sum_{j=1}^{n-1}S_{j}-\sum_{j=1}^{n-k-1}S_{j}\bigg)$
$\leq-\frac{3}{2}k(k+1)+2\big(n(n-1)-(n-k)(n-k-1)\big)$
$=4nk-\frac{7}{2}k^{2}-\frac{7}{2}k.$
And when $k$ is odd we have
$\overline{\lambda}_{n-k}$ $=\sum_{j=1}^{k}e_{n-j}-\sum_{j=1}^{k-1}\overline{\phi}_{n-j}$
$=\sum_{j=1}^{k}e_{n-j}-[(\overline{\phi}_{n-k+1}+\overline{\phi}_{n-k+2})+(\overline{\phi}_{n-k+3}+\overline{\phi}_{n-k+4})+\ldots+(\overline{\phi}_{n-2}+\overline{\phi}_{n-1})]$
$=\sum_{j=1}^{k}e_{n-j}-[(2e_{n-k+1}-S_{n-k+1})+(2e_{n-k+3}-S_{n-k+3})+\ldots+(2e_{n-2}-S_{n-2})]$
$=\sum_{j=1}^{k}(-1)^{k-j}e_{n-j}+(S_{n-k+1}+S_{n-k+3}+\ldots+S_{n-2})$
$=\sum_{j=1}^{(k-1)/2}(e_{n-2j+1}-e_{n-2j})+e_{n-k}+\sum_{j=1}^{(k-1)/2}S_{n-2j}$
$=\sum_{j=1}^{(k-1)/2}(6n-6(n-2j+1)+3-S_{n-2j+1}-2S_{n-2j})+e_{n-k}+\sum_{j=1}^{(k-1)/2}S_{n-2j}$
$=\sum_{j=1}^{(k-1)/2}(6n-6(n-2j+1)+3)+e_{n-k}-\sum_{j=1}^{(k-1)/2}(S_{n-2j+1}+S_{n-2j})$
$=\sum_{j=1}^{(k-1)/2}(12j-3)+e_{n-k}-\sum_{j=1}^{k-1}S_{n-j}$
$=\frac{3}{2}k(k-1)+e_{n-k}-\sum_{j=1}^{k-1}S_{n-j}$
$=\frac{3}{2}k(k-1)+(6(n-k)-3)n-3(n-k)^{2}-S_{n-k}-3\sum_{j=1}^{n-k-1}S_{j}-\sum_{j=1}^{k-1}S_{n-j}$
$=\frac{3}{2}k(k-1)+(6(n-k)-3)n-3(n-k)^{2}-\sum_{j=1}^{n-1}S_{j}-2\sum_{j=1}^{n-k-1}S_{j}$
$\leq\frac{3}{2}k(k-1)+(6(n-k)-3)n-3(n-k)^{2}-n(n-1)-2(n-k-1)(n-k)$
$=4nk-\frac{7}{2}k^{2}-\frac{7}{2}k.$
In both cases we have an upper bound of $\overline{\lambda}_{n-k}\leq 4nk-\frac{7}{2}k^{2}-\frac{7}{2}k$. We can combine this with our upper bound of $\lambda_{k}\leq 3nk-\frac{5}{2}k^{2}-\frac{5}{2}k$ from the proof of Theorem 3 to get
$\lambda_{k}+\overline{\lambda}_{n-k}\leq 7nk-6k^{2}-6k.$
This expression is quadratic in $k$, so it is straightforward to check that $7nk-6k^{2}-6k<\binom{n}{2}$ whenever
$k<\frac{7n-6-\sqrt{37n^{2}-72n+36}}{12}=\bigg(\frac{7-\sqrt{37}}{12}\bigg)n+O(1)\approx\frac{1}{13.08}n.$
Thus, for any $k$ that is less than approximately $\frac{n}{13.08}$, there exists a pair of points $\{p,q\}$ such that any geodesic disk with $p$ and $q$ on its boundary will contain at least $k$ points and at most nkn-k points. ∎
@@ 4.3 Sets in geodesically convex position
In the Euclidean setting, Hayward et al. [13] proved that when points are in convex position, for every disk containing two points on its boundary, there always exists a pair of points such that any disk containing that pair contains at least $\left\lceil\frac{n}{3}\right\rceil+1$ points. We generalize that result to the geodesic setting by showing that $\overline{\Pi}(n)\geq\left\lceil\frac{n}{3}\right\rceil+1$. The geodesic case poses its unique challenges, for example, it is not always possible to find an enclosing geodesic disk (i.e., containing the set $S$) with three points on the boundary. However, when no such enclosing geodesic disk exists, we show that there exists a pair of points such that every geodesic disk through them is an enclosing geodesic disk.
@@
[…]
We now consider the point $c^{\prime}$ moving along $b(y,z)$ towards the boundary of the polygon. By definition, $D(c^{\prime})$ has $y$ and $z$ on its boundary. Moreover, $D(c^{\prime})$ must enclose $S$. If $D(c^{\prime})$ is not enclosing, then there must be a point $w$ on $b(y,z)$ between $b(y,z)\cap g(x,y)$ and $c^{\prime}$ such that a third point lies on the boundary of $D(w)$, so we contradict the assumption that there is no enclosing geodesic disk through three points. The lemma follows. ∎
We now generalize the result from Hayward et al. [13] for point sets in convex position to the geodesic setting. In particular, when $S$ is a set of $n$ points in $P$ that is in geodesically convex position, then there is a pair of points such that any geodesic disk through them contains at least $\left\lceil\frac{n}{3}\right\rceil+1$ points of $S$.
@@
For any $n\geq 4$, $\overline{\Pi}(n)\geq\left\lceil\frac{n}{3}\right\rceil+1$.
@@
Otherwise, there is an enclosing disk $D$ that has three points on its boundary, namely $x,y,z$ in clockwise order along the boundary of $D$. Let $f(\triangle(x,y,z))$ be the maximum among the number of points from $S$ contained in each of the three regions $D\cap P^{+}(x,y),D\cap P^{+}(y,z)$ and $D\cap P^{+}(z,x)$. Note that, since $S$ is geodesically convex, there is no point inside the geodesic core $\triangledown$$\triangledown(x,y,z)$, and each point (other than $x,y,z$) lies in exactly one of the regions $P^{+}$. Among all enclosing disks with three points on the boundary, consider disk $D_{1}$ with points $u,v,w\in S$ on its boundary which minimizes $f(\triangle(u,v,w))$. There must be at least one region among $D_{1}\cap P^{+}(u,v),D_{1}\cap P^{+}(v,w)$ and $D_{1}\cap P^{+}(w,u)$ whose interior has at least $\left\lceil\frac{n-3}{3}\right\rceil$ points. Assume, without loss of generality, that $D_{1}\cap P^{+}(u,v)$ does, see Figure 6 (left).
@@
If the interior of $D_{1}\cap P^{-}(u,v)$ contains fewer than $\left\lceil\frac{n-3}{3}\right\rceil$ points, then we continuously change $D_{1}$ by the disks through $u,v$ obtained when moving the center $c$ of $D_{1}$ along the part of the bisector $b(u,v)$ to the right of $\ell(u,c)$, as in Figure 6. As we move $c$, we may arrive at one of two situations: Either the center $c$ hits the boundary of the polygon, or the boundary of the current disk, $D_{2}$, hits a point $w_{1}\in D_{1}\cap P^{+}(u,v)$. We show, by contradiction, that the latter case cannot occur. If it occurred, the geodesic triangle $\triangle(u,w_{1},v)$ divides $D_{2}$ into three regions $D_{2}\cap P^{+}(u,w_{1})$, $D_{2}\cap P^{+}(w_{1},v)$ and $D_{2}\cap P^{+}(v,u)$. Note that $D_{2}\cap P^{+}(v,u)=D_{2}\cap P^{-}(u,v)$ contains fewer than $\left\lceil\frac{n-3}{3}\right\rceil$ points of $S$, since it is contained in $D_{1}\cap P^{-}(u,v)$. Also, $D_{2}\cap P^{+}(u,w_{1})$ and $D_{2}\cap P^{+}(w_{1},v)$ both contain fewer points than $D_{1}\cap P^{+}(u,v)$. Thus, $f(\triangle(u,w_{1},v))<f(\triangle(u,v,w))$, and contradicting the minimality of $f(\triangle(u,v,w))$.
@@ 5 Diametral geodesic disk
If there does not exist an enclosing disk with three points on its boundary, then, by Lemma 8, there exist two points $x,y\in S$ such that every geodesic disk through them is a geodesic enclosing disk. In particular, the disk $D\left(c_{xy},\frac{\lvert g(x,y)\rvert}{2}\right)$ is an enclosing diametral disk with $x$ and $y$ as its diametral endpoints.
Suppose therefore that there exists an enclosing disk with three points on its boundary. Let $D$ be such a disk, and let $x,y,z\in S$ be the three points on the boundary of $D$, in clockwise order. If the center of the disk is inside the geodesic triangle $\triangle(x,y,z)$, we are done. Otherwise, the center is inside one of the regions $D\cap P^{+}(x,y),D\cap P^{+}(y,z)$ or $D\cap P^{+}(z,x)$. Assume, without loss of generality, it is in $D\cap P^{+}(x,y)$. Then we move the center $c$ of $D$ along $b(x,y)$ towards $c_{xy}$. If the center $c$ reaches $c_{xy}$, then $D\left(c_{xy},\frac{\lvert g(x,y)\rvert}{2}\right)$ is a diametral enclosing geodesic disk with $x$ and $y$ as its diametral endpoints. If the disk hits a third point $w\in S$, and the center of the disk is inside $\triangle(x,y,w)$, we are done. Otherwise, we repeat the process with the points $x,y,w$ as the new points $x,y,z$ until the center of the disk through the three points on the boundary of the disk lies inside the geodesic triangle formed by the three points on the boundary. This process ends because the distance from $c$ to the closest point on the geodesic $g(x,y)$ is decreasing at every iteration. ∎
@@
For any $n\geq 3$, $\Pi^{diam}(n)=\overline{\Pi}^{diam}(n)\geq\left\lceil\frac{n}{3}\right\rceil+1$.
@@
For case (i), let $a,b$ and $c$ be the three points on the boundary of the geodesic disk. We will prove that any point $x\in S$ is contained in at least one of the geodesic disks $D\left(c_{ab},\frac{\lvert g(a,b)\rvert}{2}\right)$, $D\left(c_{bc},\frac{\lvert g(b,c)\rvert}{2}\right)$ and $D\left(c_{ca},\frac{\lvert g(c,a)\rvert}{2}\right)$.
First, consider the case where $x$ is outside the geodesic triangle $\triangle(a,b,c)$. Then, $x$ belongs to one of the regions $D\cap P^{+}(a,b),D\cap P^{+}(b,c)$, and $D\cap P^{+}(a,c)$. Without loss of generality, assume $x\in D\cap P^{+}(a,b)$. By Lemma 9, the center of the geodesic disk is inside $\triangle(a,b,c)$, so it does not belong to $D\cap P^{+}(a,b)$. Then, by Lemma 1, $x$ is contained in the diametral geodesic disk $D\left(c_{ab},\frac{\lvert g(a,b)\rvert}{2}\right)$.
Now, we prove the case where $x\in\triangle(a,b,c)$. Consider the geodesic paths $g(a,x),g(b,x)$ and $g(c,x)$. One of the angles $\angle axb,\angle bxc$ or $\angle axc$ is greater than or equal to $\frac{2\pi}{3}$. Assume, without loss of generality, that $\angle bxc\geq\frac{2\pi}{3}$, see blue angle in Figure 7. Let $o$ be the midpoint of the shortest path $g(b,c)$. We also know that one of the angles $\angle bxo$ or $\angle oxc$ is greater than or equal to $\frac{\pi}{3}$. Assume, without loss of generality, that it is the angle $\angle oxc$. When $o$ is not visible from $x$, the angle $\angle oxc$ is equal to the angle $\angle bxc\geq\frac{2\pi}{3}>\frac{\pi}{2}$. Hence, by [20, Corollary $2$], $\lvert g(x,o)\rvert<\lvert g(o,c)\rvert=\lvert g(o,b)\rvert$, and $x$ is contained in the geodesic disk $D\left(c_{bc},\frac{\lvert g(b,c)\rvert}{2}\right)$. In the case of $o$ being visible from $x$, by Lemma 4, $\lvert g(o,c)\rvert\geq\min\{\lvert g(c,x)\rvert,\lvert g(x,o)\rvert\}$ and, by Lemma 5, $\lvert g(x,o)\rvert\leq\max\{\lvert g(c,x)\rvert,\lvert g(o,c)\rvert\}$, hence $\lvert g(x,o)\rvert\leq\lvert g(o,c)\rvert=\lvert g(o,b)\rvert$, and $x$ is contained in the geodesic disk $D\left(c_{bc},\frac{\lvert g(b,c)\rvert}{2}\right)$.
@@
Then, the union of the geodesic disks $D\left(c_{ab},\frac{\lvert g(a,b)\rvert}{2}\right),D\left(c_{bc},\frac{\lvert g(b,c)\rvert}{2}\right)$ and $D\left(c_{ca},\frac{\lvert g(c,a)\rvert}{2}\right)$ covers all the points in $P$. Since the points $a,b$ and $c$ are counted twice,
$\Bigg\lvert D\left(c_{ab},\frac{\lvert g(a,b)\rvert}{2}\right)\cap P\Bigg\rvert+\Bigg\lvert D\left(c_{bc},\frac{\lvert g(b,c)\rvert}{2}\right)\cap P\Bigg\rvert+\Bigg\lvert D\left(c_{ca},\frac{\lvert g(c,a)\rvert}{2}\right)\cap P\Bigg\rvert>n+3.$